OpenStudy (anonymous):

How do I solve x+5x/3=32

OpenStudy (anonymous):

balance the equation following PEMDAS http://www.coolmath.com/prealgebra/05-order-of-operations/images/order-of-operations-05-01.gif x + 5x / 3 = 32 multiply each side by 3 to get ride of the / 3 3x + 5x = 32/3 add the x values 8x = 32/3 multiply each side by 3 to get ride of the denominator again 24x = 32 divide by 24 x = 1/3